OpenStudy (anonymous):

OK on my last problem Find dy/dx in terms of x and y if 5 root x= 7Sqrt(y)

OpenStudy (anonymous):

\[\sqrt[5]{x}=7\sqrt{y}\]

OpenStudy (anonymous):

Can express with rational exponents. \(\large x^{1/5}=7y^{1/2}\) Then use power rule and implicit differentiation.

OpenStudy (anonymous):

ahh ok ty